1,1,2,3,5,8,... find the next five terms of the Fibonacci sequence
Found 2 solutions by stanbon, octogirl:
Answer by stanbon(75887)

1,1,2,3,5,8,... find the next five terms of the Fibonacci sequence
-----
5+8 = 13 is the next
8+13 = 21 is the next
13+21 = 34 is the next
21+34 = 55 is the next
34 + 55 = 89 is the next
=======================
Cheers,
Stan H.

Answer by octogirl(7)

this is very simple you just add the last two numbers
5+8=13
8+13=21
13+21=34
21+34=55
34+55=89